# DPZV: Resource Efficient ZO Optimization For Differentially Private VFL

### v1

@@ Abstract
Vertical Federated Learning (VFL) enables collaborative model training across feature-partitioned data, yet faces significant privacy risks and inefficiencies when scaling to large models. We propose DPZV, a memory-efficient Zeroth-Order(ZO) optimization framework that integrates differential privacy (DP) with vertical federated learning, addressing three critical challenges: (1) privacy vulnerabilities from gradient leakage, (2) high computation/communication costs of first-order methods, and (3) excessive memory footprint in conventional zeroth-order approaches. Our framework eliminates backpropagation through two-point gradient estimation, reducing client memory usage by 90% compared to first-order counterparts while enabling asynchronous communication. By strategically injecting Gaussian noise on the server, DPZV achieves rigorous $(\epsilon,\delta)$-DP guarantees without third-party trust assumptions. Theoretical analysis establishes a convergence rate matching centralized case under non-convex objectives. Extensive experiments on image and NLP benchmarks demonstrate that DPZV outperforms all baselines in accuracy while providing strong privacy assurances ($\epsilon\leq 10$) and requiring far fewer computation resources, establishing new state-of-the-art privacy-utility tradeoffs for resource-constrained VFL deployments.
@@ 1 Introduction
The rapid growth of data-driven applications across industries has propelled the demand for collaborative machine learning frameworks that can harness data spread across multiple organizations. In particular, vertical federated learning (VFL) has emerged as a promising paradigm to train models on data that are partitioned feature-wise among distinct parties, without directly sharing raw information across institutional boundaries [1, 2, 3]. By maintaining separate sub-models locally and exchanging only intermediate representations or gradients, VFL seeks to preserve privacy while jointly leveraging the complementary features held by different entities.
Despite its privacy-oriented design, recent studies highlight that VFL remains susceptible to data leakage through subtle channels. Two key sources of leakage have been identified: (1) feature leakage, wherein adversaries can infer sensitive inputs from the trained bottom model parameters exchanged during federated optimization, and (2) label leakage, where gradients transmitted between participants inadvertently reveal sensitive label information. These vulnerabilities underscore the urgency of integrating robust privacy-preserving techniques into VFL.
Differential privacy (DP) has emerged as a leading candidate for mitigating such vulnerabilities, as it provides strong, mathematically grounded privacy guarantees against a wide range of adversarial inferences [4, 5, 6]. However, incorporating DP mechanisms into VFL must be done carefully to avoid undue performance degradation and excessive computational overhead. The challenges are further exacerbated by the recent trend toward larger and more complex models—such as large language models—in federated scenarios, where memory and communication efficiency become paramount. Storing and propagating high-dimensional gradients or performing complex optimization steps can quickly become impractical, especially in resource-constrained settings.
An attractive alternative to conventional gradient-based optimization in VFL is Zeroth-Order (ZO) Optimization, which relies on function value evaluations rather than gradient computations [7, 8, 9]. By circumventing the need for direct gradient exchanges, ZO methods can reduce the risk of gradient-based leakage and naturally fit into memory-constrained environments. However, existing ZO-based VFL frameworks still face critical limitations. They often lack efficient communication protocols and fail to incorporate strong privacy mechanisms such as differential privacy, ultimately falling short of delivering a holistic solution that addresses privacy, scalability, and efficiency simultaneously.
In this work, we propose Differentially Private Zeroth-Order Vertical Federated Learning (DPZV), a novel ZO-based VFL framework that achieves all three objectives: privacy protection, memory and communication efficiency, and strong theoretical convergence guarantees. Our main contributions are as follows:
  • We introduce a zero-order based VFL algorithm that incorporates arbitrary levels of differential privacy at both the model and the gradient level, which is the first work in VFL to enable highly secure communication while maintaining both memory and communication efficiency. The algorithm also significantly reduces memory footprint and communication overhead associated with inter-institutional model training, making it suitable for large-scale and resource-constrained settings (Sec. 4).
  • We provide a rigorous theoretical analysis of both the convergence rate and the privacy guarantees of our method. Our convergence results establish that the proposed zero-order approach achieves the same order of convergence as the centralized counterpart, offering strong theoretical support for its scalability to large-scale systems. Furthermore, we prove that our method satisfies $(\epsilon,\delta)$-differential privacy, demonstrating its robustness in protecting privacy against potential attacks (Sec. 5 and 6).
  • We conduct extensive experiments on both classical image classification tasks and pretrained language models. We demonstrate that our algorithm outperforms state-of-the-art methods while requiring less memory to train and lower communication overhead for convergence (Sec. 1).
By jointly addressing privacy risks, memory and communication constraints, and providing formal convergence assurances, our framework represents a significant step forward for practical, secure, and scalable federated learning solutions.
@@ 2.1 Vertical Federated Learning
Vertical Federated Learning (VFL) enables collaborative training across organizations with vertically partitioned features. Early VFL frameworks focused on simple client-side models such as logistic regression and linear models [1], where clients computed local gradients for global aggregation. These methods prioritized simplicity, but lacked expressiveness for complex tasks. To address this limitation, larger client-side models like deep neural networks (DNNs) were adopted [2, 3, 10]. Studies have demonstrated that these larger client-side models can be more effective than classical methods [11]. However, larger models requires higher computation capability and VRAM cost, which calls for computation and memory efficient methods.
@@
In an effort to reduce communication cost, several communication-efficient VFL methods were proposed. One way of reducing communication overhead is by multiple local updates. [12] introduced FedBCD, a federated stochastic block coordinate descent method, which allows clients to perform multiple updates before synchronization. Similarly, Flex-VFL [3], a flexible strategy offering varying local update counts per party, constrained by a communication timeout. VIMADMM [10] adopted an ADMM-based approach to enable multiple local updates in VFL. All these methods effectively reduce the number of round-trip communications in each iteration. On the other hand, Asynchronous VFL methods (e.g., FDML [13], VAFL [2]) decouple coordination, allowing clients to update models independently. While asynchronous methods risk staleness, they significantly improve scalability in cross-silo deployments. However, in first-order methods, the backward pass typically imposes communication overhead comparable to the forward pass. In contrast, our ZO-based approach significantly reduces the cost associated with backward propagation.
@@
Privacy guarantees are critical for VFL adoption. Some VFL architectures use crypto-based privacy-preserving techniques such as Homomorphic Encryption (HE) [14], but lack formal assurances, while DP provides rigorous mathematical protection. Key DP-based methods include VAFL [2], which injects Gaussian noise into client embeddings during forward propagation to achieve Gaussian DP, and VIMADMM [10] which perturbs linear model parameters with bounded sensitivity, ensuring DP guarantees for convex settings.
2.2 Zeroth-Order Optimization in VFL
@@
Recent research has explored Zeroth-Order(ZO) Optimization within VFL to accommodate clients whose models are non-differentiable and to reduce gradient leakage. Early work like ZOO-VFL [7] adopted a naive ZO approach throughout VFL training but provided no DP guarantees. VFL-CZOFO [9] introduced a cascaded hybrid optimization method, combining zeroth-order and first-order updates, which leveraged intrinsic noise from ZO for limited privacy. However, its DP level was not adjustable, often resulting in insufficient protection.
@@
More recently, MeZO [8] proposed a memory-efficient ZO algorithm. Building upon these ideas, DPZero [15] and DPZO [16] introduced private ZO variants offering baseline privacy features. In this work, we adopt the memory-efficient design of MeZO. In addition to previous attempts to combine ZO Optimization with VFL, we integrate controllable DP into our ZO framework, which requires significantly less memory footprint.
Figure 1: An illustration of the training procedure of DPZV. The communication between the server and clients follows an asymmetric process: at each global iteration $t$, feature embeddings are uploaded once, while the server performs model updates and broadcasts ZO information to the client with DP guarantees for total $K$ iterations. This asymmetric communication update is a key design that enables memory-efficient fast convergence for our method.
3 Preliminaries
In this section, we introduce the concepts of DP and a ZO gradient estimator. Then we introduce the VFL framework, and define the objective problem.
3.1 Differential Privacy
DP is a rigorous framework designed to protect individual data while still allowing useful analysis of a dataset as a whole. We present the definition of the classical $(\epsilon,\delta)$-DP:
Definition 3.1 ($(\epsilon,\delta)$-DP).
A randomized algorithm ${\mathcal{M}}:{\mathcal{X}}^{n}\rightarrow\Theta$ is ($\epsilon,\delta$)-differentially private (DP) if for any pair of neighboring datasets $X,X^{\prime}\in{\mathcal{X}}^{n}$ that differ in the record of a single individual, and any possible event $E\subseteq\Theta$,
$\mathbb{P}[{\mathcal{M}}(X)\in E]\leq e^{\epsilon}\mathbb{P}[{\mathcal{M}}(X^{\prime})\in E]+\delta$ (1)
When considering ZO-based algorithms, model parameters are iteratively updated by an estimation of the gradient, and the updated parameter is used as the input of another iteration. This behavior involves the composition of privacy, which means in order to achieve a target privacy budget for the whole training process, we need to carefully design the privacy budget for each iteration and compute the total privacy through the composition of DP mechanisms.
3.2 Zeroth-order Optimization
Zeroth-Order Optimization has been widely discussed under both convex and non-convex conditions [17, 18, 19]. This method finds the optima of an objective function without relying on first-order information. Such properties are especially useful when first-order gradient is computationally expensive or intractable.
In the following, we introduce the two-point gradient estimator [20], which will serve as our zeroth-order gradient estimator throughout this paper.
Definition 3.2.
Let $u$ be uniformly sampled from the Euclidean sphere $\sqrt{d}{\mathbb{S}}^{d-1}$ and $\lambda>0$ be the smoothing factor. For a single datum $\xi$ sampled from the whole dataset ${\mathcal{D}}$, the two-point gradient estimator is defined as
$g(x;\xi)=\frac{f(x+\lambda u;\xi)-f(x-\lambda u;\xi)}{2\lambda}u$ (2)
In this work, we adopt the MeZO methodology [8] for our ZO Optimization, which is a memory efficient method for computing ZO information. In each communication round, every client samples a random seed $s$, which is kept private. Let the local parameters ${\bm{w}}_{m}$ and the randomly sampled direction ${\bm{u}}_{m}^{t}$ be defined as:
${\bm{w}}_{m}$ $\textstyle=[w_{m,1},w_{m,2},\dots,w_{m,d_{m}}]^{\top},$
${\bm{u}}_{m}^{t}$ $\textstyle=[u_{m,1},u_{m,2},\dots,u_{m,d_{m}}]^{\top}.$
Rather than sampling ${\bm{u}}_{m}^{t}$ as a complete vector, each entry $u_{m,j},\forall j\in[d_{m}]$, is sampled independently using the random seed $s$. The corresponding parameter $w_{m,j}$ is perturbed as follows: $w_{m,j}\leftarrow w_{m,j}\pm\lambda_{m}u_{m,j}.$ During the parameter update step, where ${\bm{u}}_{m}^{t}$ is required again, the random seed $s$ is reset, and ${\bm{u}}_{m}^{t}$ is resampled entry-wise in the same manner. This approach significantly reduces the memory footprint, halving the memory cost compared to Vanilla ZO-SGD. This reduction is particularly beneficial for client devices where memory constraints are a critical bottleneck.
3.3 Vertical Federated Learning
@@
We consider a VFL framework with one server and $M$ clients. In VFL, data is vertically partitioned between clients, with each client holding different features of the data. Suppose we have a dataset ${\mathcal{D}}$ with $D$ samples: ${\mathcal{D}}=\{\xi_{i}|i=1,2,\dots,D\}$. Each data sample $\xi_{i}$ can be partitioned into $M$ portions distributed throughout all clients, where the data sample on machine $m$ with ID $i$ is denoted as $\xi_{m,i}$, hence $\xi_{i}=[\xi_{1,i},\xi_{2,i},\ldots,\xi_{M,i}]^{\top}$. The server is numbered as machine 00 and holds the label data ${\bm{Y}}=\{y_{i}|i=1,2,\dots,D\}$.
@@ 4 Methodology
In this section, we propose DPZV, a VFL framework that combines ZO Optimization with DP guarantees and asynchronous updates. Based on (3), the objective is to collaboratively minimizing global objective function across all clients, each holds disjoint features of the same data records, and one server which holds the label data. The training procedure can be described in two iterative steps: (i) Client Update and Forward Communication: The sampled client computes local information and transmits it to the server, we define the client-server communication as forward communication. (ii) Server Update and Backward Communication: The server updates the global model and transmit global updates to the client, we define the server-client communication as backward communication.

### v2

# DPZV: Elevating the Tradeoff between Privacy and Utility in Zeroth-Order Vertical Federated Learning

@@ Abstract
Vertical Federated Learning (VFL) enables collaborative training with feature-partitioned data, yet remains vulnerable to privacy leakage through gradient transmissions. Standard differential privacy (DP) techniques like DP-SGD are difficult to apply in this setting due to VFL’s distributed nature and the high variance incurred by vector-valued noise. On the other hand, zeroth-order (ZO) optimization techniques can avoid explicit gradient exposure, but lack formal privacy guarantees. In this work, we propose DPZV, the first ZO optimization framework for VFL that achieves tunable DP with performance guarantees. DPZV overcomes the aforementioned limitations by injecting low-variance scalar noise at the server, enabling controllable privacy with reduced memory overhead. We conduct a comprehensive theoretical analysis showing that DPZV matches the convergence rate of first order (FO) optimization methods while satisfying formal $(\epsilon,\delta)$-DP guarantees. Experiments on image and language benchmarks demonstrate that DPZV outperforms several baselines in terms of achieved accuracy under a wide range of privacy constraints ($\epsilon\leq 10$), thereby elevating the privacy-utility tradeoff in VFL.
@@ 1 Introduction
Vertical Federated Learning (VFL) has emerged as a compelling paradigm for training models on data that are partitioned feature-wise among distinct parties [15, 4, 3]. It seeks to preserve privacy by maintaining separate submodels locally and exchanging only intermediate representations or gradients, without exposing their raw data. Despite its privacy-aware architecture, recent studies have demonstrated that VFL remains vulnerable to feature leakage [19, 36] and label leakage [13, 41], where adversaries can infer sensitive input attribute or labels from the exchanged model updates or intermediate gradients. These findings highlight a critical need for integrating stronger, principled privacy-preserving mechanisms into the VFL framework.
One direction for mitigating privacy risks in VFL is through replacing SGD-based updates with zeroth-order (ZO) optimization[28]. By avoiding the transmission of intermediate gradients, ZO naturally reduces the risk of label leakage[39]. However, the privacy level of ZO-based methods are not directly adjustable, which constrains their applicability in settings that demand adaptive privacy guarantees aligned with task-specific requirements. Moreover, ZO remains susceptible to feature leakage, since sensitive information may still be inferred from local model parameters.
Importantly, current ZO-based VFL algorithms fail to provide mathematically rigorous differential privacy (DP) [10] guarantees, unlike their SGD counterparts. The DP mechanism is desirable as it allows for adjustable privacy levels under varying privacy budgets, enabling flexible trade-offs between model performance and privacy[11]. It also offers protection against a broad class of inference attacks  [40, 31, 22]. As a result, in this work, we ask the following research question:
How can we incorporate a controllable differential privacy mechanism into zeroth-order VFL which maintains convergence guarantees while elevating the privacy-utility tradeoff?
Key Challenges. Integrating differential privacy into ZO-based VFL poses several fundamental challenges. In particular, existing strategies for achieving DP within the VFL framework typically involve injecting vector-valued noise into either (i) forward embeddings or (ii) backward gradients[4, 35]. While approach (i) can provide formal privacy guarantees, it requires adding noise with the same shape as the transmitted embeddings, which are typically high-dimensional. This high-dimensional noise introduces substantial variance amplification[1]. Moreover, ZO methods inherently produce noisy gradient estimates[14]: when combined with the additional variance introduced by DP noise, the compounded effect can lead to substantial performance degradation. This situation can be further exacerbated when the privacy budget is tight. On the other hand, approach (ii) is incompatible with ZO methods, as ZO does not involve transmitting backward gradients[12].
Contributions. To address the above challenges, we propose Differentially Private Zeroth-Order VFL (DPZV), a novel ZO-based VFL framework that achieves a controllable differential privacy level with convergence guarantees. By injecting scalar-valued noise rather than high-dimensional vector noise, DPZV mitigates variance amplification, thereby enhancing the tradeoff between privacy and model performance in VFL. Our main contributions can be summarized as follows:
  • DPZV is the first ZO optimization framework for vertical federated learning (VFL) that enables tunable differential privacy. By injecting calibrated scalar-valued noise instead of high-dimensional vector noise, DPZV provides privacy guarantees while preserving model accuracy. Moreover, the use of ZO optimization substantially reduces memory overhead, making DPZV particularly suitable for privacy-critical and resource-constrained environments (Sec. 4).
  • We provide a rigorous theoretical analysis establishing both convergence and privacy guarantees for DPZV. Specifically, we show that DPZV achieves a convergence rate on the same order as first-order DP-SGD, despite using only ZO estimators. This result indicates that our DPZV framework attains privacy without sacrificing training efficiency. Furthermore, we prove that our method satisfies $(\epsilon,\delta)$-DP, confirming its ability to provide an adjustable privacy control mechanism (Sec. 5 and 6).
  • Through experiments on image and language benchmarks, we demonstrate that DPZV obtains robust convergence performance under strict privacy budgets. While other DP baselines experience a steep performance degradation as the privacy level increases, DPZV consistently maintains high accuracy across all tasks, underscoring an elevated tradeoff between privacy and utility (Sec. 7).
@@ 2 Related Work
Vertical Federated Learning. Vertical Federated Learning (VFL) enables collaborative training across organizations with vertically partitioned features. Early VFL frameworks focused on simple client-side models such as logistic regression and linear models [15], where clients computed local gradients for global aggregation. These methods prioritized simplicity, but lacked expressiveness for complex tasks. To address this limitation, larger client-side models like deep neural networks (DNNs) were adopted [4, 3, 35]. Studies have demonstrated that these larger client-side models can be more effective than classical methods [23]. However, larger models require higher processing capability and VRAM costs, which calls for computation and memory efficient methods.
@@
In an effort to reduce communication cost, several communication-efficient VFL methods have been proposed. One popular mechanism for reducing communication overhead is by allowing for multiple local updates in-between aggregations. In this respect, [24] introduced FedBCD, a federated stochastic block coordinate descent method, which allows clients to perform multiple gradient iterations before synchronization. Similarly, Flex-VFL [3] proposed a flexible strategy offering varying local update counts per party, constrained by a communication timeout. VIMADMM [35] adopted an ADMM-based approach to enable multiple local updates in VFL. All these methods effectively reduce the number of round-trip communications in each iteration. On the other hand, Asynchronous VFL methods (e.g., FDML [17], VAFL [4]) decouple coordination, allowing clients to update models independently. While asynchronous methods risk staleness, they significantly improve scalability in cross-silo deployments. However, in first-order methods, the backward pass through neural networks typically imposes communication overhead comparable to the forward pass. In contrast, our ZO-based approach significantly reduces the cost associated with backward propagation.
@@
Privacy guarantees are critical for VFL adoption. Some VFL architectures use crypto-based privacy-preserving techniques such as Homomorphic Encryption (HE) [6], but lack formal assurances, while DP provides rigorous mathematical protection. Key DP-based methods include VAFL [4], which injects Gaussian noise into client embeddings during forward propagation to achieve Gaussian DP, and VIMADMM [35], which perturbs linear model parameters with bounded sensitivity, ensuring DP guarantees for convex settings. Our work overcomes challenges in developing such methods for the ZO setting.
@@
Zeroth-Order Optimization. Recent research has explored Zeroth-Order (ZO) optimization within VFL to accommodate resource-constrained clients with non-differentiable models and to reduce gradient leakage. Early work like ZOO-VFL [39] adopted a naive ZO approach throughout VFL training but provided no DP guarantees. VFL-CZOFO [33] introduced a cascaded hybrid optimization method, combining zeroth-order and first-order updates, which leveraged intrinsic noise from ZO for limited privacy. However, its DP level was not adjustable, often resulting in insufficient protection.
@@
More recently, MeZO [26] proposed a memory-efficient ZO algorithm. Building upon these ideas, DPZero [38] and DPZO [32] introduced private ZO variants offering baseline privacy features. However, these methods are designed for centralized settings and cannot be directly extended to the VFL paradigm. In this work, we incorporate the memory-efficient design of MeZO into a VFL framework. Extending beyond previous efforts to combine ZO optimization with VFL, we further integrate a controllable differential privacy mechanism, achieving an elevated trade-off between privacy and model performance.
Figure 1: Overview of the training procedure in DPZV. Each client perturbs its local model parameters in two random directions to generate a pair of embeddings, which are then transmitted to the server. The server computes the corresponding function evaluations and applies an elementwise difference to approximate the zeroth-order (ZO) gradient. To ensure differential privacy, scalar-valued Gaussian noise is injected into the aggregated ZO estimate. Unlike traditional vector-valued noise in standard DP algorithms, scalar noise is significantly smaller in norm, thereby preserving model utility even under stringent privacy budgets.
3 Background and Motivation
@@
System Model for VFL. We consider a VFL framework with one server and $M$ clients. In VFL, data is vertically partitioned between clients, with each client holding different features of the data. Suppose we have a dataset ${\mathcal{D}}$ with $D$ samples: ${\mathcal{D}}=\{\xi_{i}|i=1,2,\dots,D\}$. Each data sample $\xi_{i}$ can be partitioned into $M$ portions distributed throughout all clients, where the data sample on machine $m$ with ID $i$ is denoted as $\xi_{m,i}$, hence $\xi_{i}=[\xi_{1,i},\xi_{2,i},\ldots,\xi_{M,i}]^{\top}$. The server is numbered as machine 00 and holds the label data ${\bm{Y}}=\{y_{i}|i=1,2,\dots,D\}$.
@@
Memory-Efficient Zeroth-Order Optimizer. We introduce the two-point gradient estimator [30], which will serve as our zeroth-order gradient estimator throughout this paper. Let $u$ be uniformly sampled from the Euclidean sphere $\sqrt{d}{\mathbb{S}}^{d-1}$ and $\lambda>0$ be the smoothing factor. For a single datum $\xi$ sampled from the whole dataset ${\mathcal{D}}$, the two-point gradient estimator is defined as
$\textstyle g(x;\xi)=\frac{f(x+\lambda u;\xi)-f(x-\lambda u;\xi)}{2\lambda}u$ (3)
To further reduce the memory overhead on client, we adopt the MeZO methodology [26] for our ZO Optimization, which requires only the same memory as the model itself for training. We investigate the memory reduction in Appendix  F.
Differential Privacy and DP-SGD. Differential Privacy (DP) provides a principled framework for protecting individual data in statistical analysis and machine learning. Formally, we adopt the standard $(\epsilon,\delta)$-DP definition:
Definition 3.1 ($(\epsilon,\delta)$-Differential Privacy).
A randomized algorithm ${\mathcal{M}}:{\mathcal{X}}^{n}\rightarrow\Theta$ is said to satisfy $(\epsilon,\delta)$-differential privacy if for any neighboring datasets $X,X^{\prime}\in{\mathcal{X}}^{n}$ that differ by only one individual’s data, and for any subset of outputs $E\subseteq\Theta$, we have
$\mathbb{P}[{\mathcal{M}}(X)\in E]\leq e^{\epsilon}\mathbb{P}[{\mathcal{M}}(X^{\prime})\in E]+\delta.$ (4)
One of the most widely adopted private training algorithms is Differentially Private Stochastic Gradient Descent (DP-SGD), which operates by clipping an intermediate result and injecting Gaussian noise during each model update. Specifically, given a minibatch $B_{t}$ at iteration $t$, DP-SGD performs the following steps: 1) Clip individual gradients: $\bar{g}_{i}=\nabla\ell(\theta_{t},x_{i}),\quad\tilde{g}_{i}=\bar{g}_{i}/\max\left(1,\frac{\|\bar{g}_{i}\|_{2}}{C}\right)$. 2) Add Gaussian noise: $\tilde{g}_{t}=\frac{1}{|B_{t}|}\left(\sum_{i\in B_{t}}\tilde{g}_{i}+\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I})\right)$. 3) Update model: $\theta_{t+1}=\theta_{t}-\eta_{t}\tilde{g}_{t}$.
Challenges and Motivation. While DP-SGD is effective in centralized or horizontally federated setups, it becomes more challenging in VFL, particularly when ZO optimization is employed. First, due to the distributed nature of VFL, differential privacy must be enforced before the server communicates with clients to protect against potential inference attacks by malicious client parties. In both cases, the noise must match the dimensionality of the protected vector (embedding or gradient), often resulting in high-dimensional noise. This situation is further exacerbated by the second factor: in ZO methods, gradients are approximated through multiple forward evaluations, making them inherently noisy. Adding a high-dimensional noise vector as in standard DP-SGD further amplifies this inaccuracy.
A notable benifit of ZO optimization is its communication efficiency in the backward pass, where only a scalar value instead of high-dimensional gradients are transmitted. This enables a more efficient privacy mechanism—injecting scalar noise rather than high-dimensional noise. Motivated by this observation, we develop Differentially Private Zeroth-Order Vertical Federated Learning (DPZV), a framework that achieves tunable differential privacy through calibrated scalar noise injection. This design allows for a favorable privacy-utility trade-off in ZO-based VFL, particularly under tight privacy budgets.
@@ 4 Methodology
In this section, we describe the overall training procedure of DPZV. Based on (1), the objective is to collaboratively minimizing global objective function across all clients, each holds disjoint features of the same data records, and one server which holds the label data. The training procedure can be described in two iterative steps: (i) Client Update and Forward Communication: The sampled client computes local information and transmits it to the server, we define the client-server communication as forward communication. (ii) Server Update and Backward Communication: The server updates the global model and transmit global updates to the client, we define the server-client communication as backward communication.
Input : Data ${\mathcal{D}}$, batch size $B$, learning rate $\eta_{m}$, total iteration $T$, smoothing parameter $\lambda_{m}>0$, clipping threshold $C>0$, privacy parameter $\sigma_{dp}$
Output : Parameter $w_{0},w_{m}$ for all parties $m\in[M]$
1 Initialize $w_{0},w_{m}$ and set $t,t_{m}=0$ for all parties
2 for $t=1,\ldots,T$ do
3       Sample ready-to-update client $m\in\{1,\ldots,M\}$.
4       Client $m$ samples a mini-batch ${\mathcal{B}}_{m}^{t_{m}}$ with corresponding IDs ${\mathcal{I}}_{m}^{t_{m}}$.
5       Client $m$ computes perturbed local embeddings $\{h_{m,i}^{t_{m}+},h_{m,i}^{t_{m}-}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}$ according to (6).
6       (Forward) Client transmits embeddings to server.
7       Server computes $\Delta_{m}^{t}$ according to (7) and (8), and updates via ZO Optimization.
8
9      (Backward) Client receives $\Delta_{m}^{t}$.
10
11      Client performs local update using (9).
12
13Client update local time stamp $t_{m}=t$.
14